Given a distribution of weight (162.9 lbs with standard deviation of 65.6 lbs), what sample size would be within 10 lbs of the population mean 90% of the time?

Use Matlab to find the sample size.
Answer #1

Standard Deviation ,   σ =    65.6                  
sampling error ,    E =   10                  
Confidence Level ,   CL=   90%                  
                          
alpha =   1-CL =   10%                  
Z value =    Zα/2 =    1.645   [excel formula =normsinv(α/2)]              
                          
Sample Size,n = (Z * σ / E )² = (   1.645   *   65.6   /   10   ) ² =   116.429
                          
                          
So,Sample Size needed=       117